OpenStudy (anonymous):

10(2^x) =640, what is the value of x?

OpenStudy (anonymous):

2^x = 64 x = 6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

and also could you help me with one more

OpenStudy (anonymous):

\[27^{x} = 1/9\] what is the value of x?

OpenStudy (anonymous):

-2/3 = x

OpenStudy (anonymous):

Press good answer if my answer is correct

OpenStudy (anonymous):

hmm.. i'm not so sure, about that one so could explain

OpenStudy (anonymous):

Log (1/9) / log 27 = -2/3